### v1

# Magnon Correlation Enables Spin Injection, Dephasing, and Transport in Canted Antiferromagnets

Abstract
Thermal and electrical injection and transport of magnon spins in magnetic insulators is conventionally understood by the non-equilibrium population of magnons. However, this view is challenged by several recent experiments in noncollinear antiferromagnets, which urge a thorough theoretical investigation at the fundamental level. We find that the magnon spin in antiferromagnets is described by a matrix, so even when the diagonal terms—spins carried by population—vanish, the off-diagonal correlations transmit magnon spins. Our quantum theory shows that a net spin-flip of electrons in adjacent conductors creates quantum coherence between magnon states, which transports magnon spins in canted antiferromagnets, even without a definite phase difference between magnon modes in the incoherent process. It reveals that the pumped magnon correlation is not conserved due to an intrinsic spin torque, which causes dephasing and strong spatial spin oscillations during transport; both are enhanced by magnetic fields. Spin transfer to proximity conductors can cause extrinsic dephasing, which suppresses spin oscillations and thereby gates spin transport.
Introduction.—Antiferromagnets (AFMs) hold significant promise for spintronic applications due to their terahertz-frequency dynamics, immunity to external magnetic disorder, and ultrafast responses [1, 2, 3, 4, 5]. Spin angular momentum in collinear easy-axis AFMs is carried by circularly polarized magnon eigenmodes, which shows low dissipation in insulators with efficient long-range spin transport [6, 8, 9, 7, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 21, 22, 20, 23, 24]. In contrast, magnon modes in easy-plane AFMs are linearly polarized and are believed to carry no spins [3, 25, 26, 27] and the modes in canted AFMs have no spin projection along the Néel vector; both appear to be less likely to support spin currents [3, 16, 28, 29]. Experiments on canted easy-plane hematite ($\alpha$-Fe2O3) deny this intuition by observing efficient long-range magnon spin transport and a field-tunable Hanle effect [30, 31, 32]. The observation is interpreted as an interference of two linearly polarized modes, occasionally mimicking a circular polarization or pseudospin [5, 33, 34, 35].
This pseudospin description so far is classical in the sense that two classical waves with a time-dependent phase difference interfere [36]. On the other hand, the thermal and electrical spin injection via the coupling between electrons and magnons across interfaces is a highly incoherent process since no definite initial phase difference between two waves is generated [6, 7, 14, 15, 17, 16, 37, 38, 39, 40, 41], different from microwave excitation [36] or spin pumping from AFMs to metals [21, 22, 20, 42, 43, 44, 45]. This essentially raises a fundamental question: How is observable macroscopic quantum coherence established between different magnon branches in the incoherent thermal/electrical injection, and how do they dephase to the environment?
Figure 1: (a) Description of magnon spin in AFMs in terms of a matrix. (b) A net spin-flip of an electron in normal metals creates the magnon correlation between two bands in AFMs. An optional configuration to pump magnon correlation is to use canted AFMs with Néel vector along $\hat{\bf z}$ driven by a spin accumulation $\parallel\hat{\bf z}$ created by a charge current $I$ along $\hat{\bf y}$ via the spin Hall effect.
In this Letter, we demonstrate in the quantum description that the pseudospin in the classical description is a “real spin” carried by magnons by showing that the magnon spin is described by a matrix. When the spin carried by magnon population vanishes due to specific interactions, as shown in Fig. 1(a), the spin is solely carried by correlation ${\cal A}$, represented by the wavy line between two magnon bands $\hbar\omega_{1}$ and $\hbar\omega_{2}$ in Fig. 1(b). By developing a quantum kinetic theory for electrical spin injection, dephasing, and transport in noncollinear AFMs, we show that a net spin-flip process of electrons creates definite magnon correlations, illustrated in Fig. 1(b) with canted AFMs as a prototypical example. The injected magnon correlation experiences dephasing due to free-induction decay, governed by the inhomogeneous broadening [48, 46, 47] of oscillation frequencies $|\omega_{1}({\bf k})-\omega_{2}({\bf k})|$ of wave vectors ${\bf k}$, and backaction to spin-flips of electrons. This then provides two routes to control the magnon-correlation transport away from the injector: i) the magnetic field causes the spin spatial oscillation in the transport that is enhanced by the applied magnetic field; ii) a proximity conductor brakes the transport, suppresses spatial oscillation, and thereby acts as a gate. Inducing quantum coherence in incoherent processes may lead to crucial applications and new tunability in future devices.
Spin carried by magnon correlations.—We start from a general Heisenberg Hamiltonian for AFMs $\hat{H}_{m}=(1/2)\sum_{l,l^{\prime}=1}^{\cal N}\sum_{i,j=1}^{N}\sum_{\alpha,\beta}{J}_{ij}^{\alpha\beta}(l,l^{\prime})\hat{S}_{li}^{\alpha}\hat{S}_{l^{\prime}j}^{\beta}$ with noncollinear spins $\hat{\bf S}_{li}$ of the $l$-th ion in the $i$-th magnetic unit cell coupled via the coupling constant $J_{ij}^{\alpha\beta}$. The magnon $\hat{\Psi}_{\bf k}=(\hat{b}_{\xi,{\bf k}},\hat{b}^{\dagger}_{\xi,{\bf-k}})^{T}$ of modes $\xi=\{{1,\cdots,{\cal N}}\}$ and wave vector ${\bf k}$ is found by Bogoliubov transformation with matrix ${\cal T}_{\bf k}$ [49], see Supplemental Material (SM) [50]. The $\alpha$-component of the total spin $\hat{S}_{\alpha}=(1/2)\sum_{\bf k}\hat{\Psi}^{\dagger}_{\bf k}{\cal S}_{\alpha}({\bf k})\hat{\Psi}_{\bf k}$ carried by magnons defines the magnon spin via a matrix
${\cal S}_{\alpha}({\bf k})=-{\cal T}^{\dagger}_{\bf k}\left[I_{{\cal N}\times{\cal N}}\otimes{\rm diag}(\eta^{\alpha}_{1},\cdots,\eta^{\alpha}_{\cal N})\right]{\cal T}_{\bf k},$ (1)
in which $\eta^{\alpha}_{l=\{1,\cdots,{\cal N}\}}$ is the projection of the $l$-th spin ${\bf S}_{l}$ along the lab $\alpha$-axis. The diagonal elements of ${\cal S}_{\alpha}({\bf k})$ are the magnon spin/polarization defined in the classical description [27, 51]. The off-diagonal elements are the magnon correlations; its physical meaning may be fully appreciated in the spin current $\hat{\bf J}_{\alpha}({\bf r})$ and torque $\hat{T}_{\alpha}({\bf r})$ with $\alpha$ denoting the spin polarization. The rate of change of spin density $\hat{S}_{\alpha}({\bf r})$ coverts to the spin current according to $\partial_{t}\hat{S}_{\alpha}({\bf r})=(1/i\hbar)[\hat{S}_{\alpha}({\bf r}),\hat{H}_{m}]=-\nabla\cdot\hat{\bf J}_{\alpha}({\bf r})+\hat{T}_{\alpha}({\bf r})$. When flowing uniformly, the spin current and torque read [52, 53, 50]
$\hat{\bf J}_{0\alpha}$ $=\frac{1}{4V}\sum_{{\bf k}}\hat{\Psi}^{\dagger}_{{\bf k}}\left({\cal S}_{\alpha}({\bf k})\sigma_{3}{\boldsymbol{\cal V}}_{\bf k}+{\boldsymbol{\cal V}}_{\bf k}\sigma_{3}{\cal S}_{\alpha}({\bf k})\right)\hat{\Psi}_{{\bf k}},$ (2a)
$\hat{T}_{0\alpha}$ $=\frac{1}{2V}\frac{1}{i\hbar}\sum_{{\bf k}}\hat{\Psi}^{\dagger}_{{\bf k}}\left({\cal S}_{\alpha}({\bf k})\sigma_{3}{\cal E}_{\bf k}-{\cal E}_{\bf k}\sigma_{3}{\cal S}_{\alpha}({\bf k})\right)\hat{\Psi}_{{\bf k}},$ (2b)
expressed by the spin matrix (1), mode energies ${\cal E}_{\bf k}={\rm diag}(\hbar\omega_{\xi,\bf k},\hbar\omega_{\xi,\bf-k})$, metric $\sigma_{3}={\rm diag}(I_{{\cal N}\times{\cal N}},-I_{{\cal N}\times{\cal N}})$, and magnon group velocities ${\boldsymbol{\cal V}}_{{\bf k}}=(1/\hbar)\partial{\cal E}_{\bf k}/\partial{\bf k}$.
[…]
Figure 2: (a) Two-sublattice model of AFM. The dashed square represents a unit cell with a lattice constant $a$. (b) and (c) Polarizations ${\bf m}$ by green arrows of the anti-phase mode “$1$” and in-phase mode “$2$” with the Néel vector ${\bf n}$ of mode “1” (“2”) oscillating along the $\hat{\bf y}~{}(\hat{\bf x})$-axis, indicated by orange arrows. (d) Band structure of two modes. Inset: average frequency difference $|\delta\omega|=|\omega_{1}-\omega_{2}|$ over $|{\bf k}|\in[0,0.5]$ nm-1.
According to the spin matrix (1), the total spin operator along the Néel vector
$\hat{S}_{z}=\sum_{{\bf k}}\left({\cal A}_{\bf k}\hat{b}_{1,{\bf k}}^{\dagger}\hat{b}_{2,{\bf k}}+{\cal B}_{\bf k}\hat{b}_{1,{\bf k}}^{\dagger}\hat{b}^{\dagger}_{2,{-\bf k}}\right)+{\rm H.c.},$ (3)
with expressions for ${\cal A}_{\bf k}$ and ${\cal B}_{\bf k}$ with arbitrary canted angles $\theta$ given in SM [50]. The population of both magnon modes carries no spin along the Néel vector; the spin is solely carried by the correlations between the two modes. A solution ${\cal A}_{\bf k}\approx 1$ and ${\cal B}_{\bf k}\approx 0$ when the canted angle is small with $|\gamma|\hbar\mu_{0}H\ll 8JS$ is sufficient to analyze the mechanism. The spin current and torque for spin polarization along ${\bf n}$ are also solely carried by magnon correlations. Indeed, according to Eqs. (2a) and (2b),
$\hat{J}_{0z}$ $\approx\frac{1}{2V}\sum_{{\bf k}}{\cal A}_{\bf k}({\bf v}_{1,{\bf k}}+{\bf v}_{2,{\bf k}})\hat{b}_{1,{\bf k}}^{\dagger}\hat{b}_{2,{\bf k}}+{\rm H.c.},$ (4a)
$\hat{T}_{0z}$ $\approx\frac{1}{V}\sum_{{\bf k}}i{\cal A}_{\bf k}(\omega_{1,{\bf k}}-\omega_{2,{\bf k}})\hat{b}_{1,{\bf k}}^{\dagger}\hat{b}_{2,{\bf k}}+{\rm H.c.},$ (4b)
without contribution from the magnon population. Due to the breaking of $U(1)$ symmetry in canted AFMs, $S_{z}$ is no longer conserved by a spin torque $\hat{T}_{0z}$.
Electrical injection of magnon correlation.—We examine the electrical spin injection from heavy metals to canted AFMs in a bilayer structure as illustrated in Fig. 1(b). When a charge current is applied along the $\hat{\bf y}$-direction in the metals, the injected spin current along $\hat{\bf x}$ generates a spin accumulation $\mu_{\uparrow}\neq\mu_{\downarrow}$ polarized along the Néel vector $\hat{\bf z}$ at the interface through the spin Hall effect. This accumulated spin of electrons then excites magnon correlations in the underlying AFM as in Fig. 1(b) via the $s$-$d$ exchange interaction between local spins and conduction electrons at the interface, governed by the Hamiltonian $\hat{H}_{\rm int}=\sum_{i\in{\rm int}}^{N_{\rm int}}\sum_{l=1}^{\cal N}{\cal J}_{l}\hat{\bf s}_{i}\cdot\hat{\bf S}_{li}$ [56, 57, 58], where the summation is performed over the magnetic unit cells “$i$”, with $N_{\rm int}$ being their total number in the interfacial plane, and the spin sublattices “$l$”. ${\cal J}_{l}$ represents the local exchange coupling, which typically depends on the AFM sublattices. Equal or unequal coupling with different sublattices corresponds to compensated and uncompensated interfaces, as observed experimentally [59, 38, 45].
For magnons in the two-sublattice model with a lattice constant $a$ [Fig. 2(a)] interacting with the electrons in a heavy metal of thickness $d$ [50, 56, 57, 58],
$\hat{H}_{\rm int}$ $=\frac{a}{d}\sqrt{\frac{S}{2N_{\rm int}}}\sum_{{\bf k}{\bf k}^{\prime}{\bf q}}\sum_{\zeta=1,2}\sum_{\alpha\beta}{\boldsymbol{\sigma}}_{\alpha\beta}\cdot{\boldsymbol{\chi}}^{(\zeta)}_{\bf q}$
×$\times\hat{c}_{{{\bf k}^{\prime}},\alpha}^{\dagger}\hat{c}_{{{\bf k}},\beta}\hat{b}_{\zeta,{\bf q}}\delta_{{{\bf k}}+{\bf q},{\bf k}^{\prime}}+{\rm H.c.},$ (5)
in which ${\boldsymbol{\sigma}}=(\sigma_{x},\sigma_{y},\sigma_{z})$ is a vector of Pauli matrices, $\hat{c}_{{\bf k},\beta}$ represents the annihilation operator of electrons, and the vectors with the superscript indicating the modes
$\boldsymbol{\chi}^{(1)}_{\bf q}$ $={\cal J}_{1}\delta{\bf S}^{(1)}_{1,{\bf q}}+{\cal J}_{2}\delta{\bf S}^{(1)}_{2,{\bf q}},$
$\boldsymbol{\chi}^{(2)}_{\bf q}$ $={\cal J}_{1}\delta{\bf S}^{(2)}_{1,{\bf q}}+{\cal J}_{2}\delta{\bf S}^{(2)}_{2,{\bf q}},$ (6)
are governed by exchange couplings ${\cal J}_{1,2}$ and polarizations of sublattice spins $\delta{\bf S}_{1,{\bf q}}^{(1)}=(\sqrt{2}/2)(-B_{1,\bf q}\cos\theta\hat{\bf x}+iA_{1,\bf q}\hat{\bf y}+B_{1,\bf q}\sin\theta\hat{\bf z})$ and $\delta{\bf S}_{2,{\bf q}}^{(1)}=(\sqrt{2}/2)(-B_{1,\bf q}\cos\theta\hat{\bf x}-iA_{1,\bf q}\hat{\bf y}-B_{1,\bf q}\sin\theta\hat{\bf z})$ in mode “$1$” and $\delta{\bf S}_{1,{\bf q}}^{(2)}=(\sqrt{2}/2)(A_{2,{\bf q}}\cos\theta\hat{\bf x}-iB_{2,{\bf q}}\hat{\bf y}-A_{2,{\bf q}}\sin\theta\hat{\bf z})$ and $\delta{\bf S}_{2,{\bf q}}^{(2)}=(\sqrt{2}/2)(-A_{2,{\bf q}}\cos\theta\hat{\bf x}-iB_{2,{\bf q}}\hat{\bf y}-A_{2,{\bf q}}\sin\theta\hat{\bf z})$ in mode “$2$”, as shown in Fig. 2(b) and (c). When the exchange interaction $8JS\gg\hbar\omega_{1,2}({\bf q})$ dominates over the other energies, $A_{1,2}({\bf q})\approx\sqrt{8JS/\hbar\omega_{1,2}({\bf q})}$ and $B_{1,2}({\bf q})\approx 1/A_{1,2}({\bf q})$, and, thereby, $A_{1,2}({\bf q})\gg B_{1,2}({\bf q})$. When ${\cal J}_{1}={\cal J}_{2}$, $|\boldsymbol{\chi}^{(1),(2)}_{\bf q}|\propto B_{1,2}({\bf q})$; when ${\cal J}_{1}\gg{\cal J}_{2}$ or ${\cal J}_{2}\gg{\cal J}_{1}$, $|\boldsymbol{\chi}^{(1),(2)}_{\bf q}|\propto A_{1,2}({\bf q})$, so the electron-magnon coupling is enhanced when ${\cal J}_{1}\neq{\cal J}_{2}$.
$|\boldsymbol{\sigma}_{\alpha\beta}\cdot\boldsymbol{\chi}^{(\zeta)}_{\bf q}|^{2}$ is responsible for the excitation of magnon population in the $\zeta$-mode; $(\boldsymbol{\sigma}_{\alpha\beta}\cdot\boldsymbol{\chi}^{(1)}_{\bf q})(\boldsymbol{\sigma}_{\beta\alpha}\cdot\boldsymbol{\chi}^{(2)\ast}_{\bf q})$ denotes the mixing excitation of both modes in a net spin-flip process, which creates a magnon correction as in Fig. 1(b). The latter process vanishes in collinear AFMs. Since the magnon population carries no spin along $\hat{\bf z}$, the electron spin is not conserved in the spin-flip scattering (5) since $\boldsymbol{\chi}^{(1,2)}_{\bf q}\nparallel\hat{\bf z}$. When the canted angle $\theta$ is small, $\chi_{z}\rightarrow 0$ such that the interfacial spin-conserving scattering $|\boldsymbol{\sigma}_{\alpha\alpha}\cdot\boldsymbol{\chi}^{(\zeta)}_{\bf q}|^{2}\rightarrow 0$. Consequently, the excitation of magnon correlation is dominated by electron spin-flip scattering.
[…]
$\partial_{t}{\rho}_{\bf q}|_{\rm inj}$ $=-\frac{\pi}{\hbar}\frac{S}{2N_{\rm int}}\frac{a^{2}}{d^{2}}\sum_{{\bf k},{\bf k}^{\prime}}\sum_{\alpha\beta}\left[{\rho}_{\bf q}-\sigma_{3}N_{0}(\varepsilon_{{\bf k}^{\prime},\alpha}-\varepsilon_{{\bf k},\beta})\right]$
×$\times(f_{{\bf k},\beta}-f_{{\bf k}^{\prime},\alpha}){\cal M}_{\alpha\beta}({\bf k},{\bf k}^{\prime})\delta_{{\bf k}+{\bf q},{\bf k}^{\prime}}+{\rm H.c.},$ (7)
where $N_{0}(\varepsilon)=\{\exp[{\varepsilon}/(k_{B}T)]-1\}^{-1}$ is the Bose-Einstein distribution at temperature $T$ and the scattering matrix
${\cal M}_{\alpha\beta}({\bf k},{\bf k^{\prime}})=\begin{pmatrix}{{\boldsymbol{\sigma}}_{\alpha\beta}\cdot{\boldsymbol{\chi}}^{(1)}_{{\bf q}}}\delta(\varepsilon_{{\bf k}}-\varepsilon_{{\bf k}^{\prime}}+\hbar\omega_{1,{\bf q}})\\ {{\boldsymbol{\sigma}}_{\alpha\beta}\cdot{\boldsymbol{\chi}}^{(2)}_{{\bf q}}}\delta(\varepsilon_{{\bf k}}-\varepsilon_{{\bf k}^{\prime}}+\hbar\omega_{2,{\bf q}})\\ {{\boldsymbol{\sigma}}_{\alpha\beta}\cdot{\boldsymbol{\chi}}_{-{\bf q}}^{(1)\ast}}\delta(\varepsilon_{{\bf k}}-\varepsilon_{{\bf k}^{\prime}}-\hbar\omega_{1,-{\bf q}})\\ {{\boldsymbol{\sigma}}_{\alpha\beta}\cdot{\boldsymbol{\chi}}_{-{\bf q}}^{(2)\ast}}\delta(\varepsilon_{{\bf k}}-\varepsilon_{{\bf k^{\prime}}}-\hbar\omega_{2,-{\bf q}})\end{pmatrix}$
×$\times\begin{pmatrix}{\boldsymbol{\sigma}}_{\beta\alpha}\cdot{\boldsymbol{\chi}}_{{\bf q}}^{(1)\ast},&{\boldsymbol{\sigma}}_{\beta\alpha}\cdot{\boldsymbol{\chi}}_{{\bf q}}^{(2)\ast},&-{\boldsymbol{\sigma}}_{\beta\alpha}\cdot{\boldsymbol{\chi}}^{(1)}_{-{\bf q}},&-{\boldsymbol{\sigma}}_{\beta\alpha}\cdot{\boldsymbol{\chi}}^{(2)}_{-{\bf q}}\end{pmatrix}.$
The electron population $f(\varepsilon_{{\bf k},\beta})=f_{0}(\varepsilon_{{\bf k}}-\varepsilon_{F})-\delta\mu_{\beta}{\partial f_{0}}/{\partial\varepsilon_{{\bf k}}}$ deviates from the equilibrium $f_{0}(\varepsilon_{{\bf k}}-\varepsilon_{F})=\{{\rm exp}[(\varepsilon_{{\bf k}}-\varepsilon_{F})/(k_{B}T)]+1\}^{-1}$ by a spin accumulation $\delta\mu_{\beta}$ that is much smaller than the Fermi energy $\varepsilon_{F}$, which exchanges the electron spin with the magnon population and correlation. Magnon population and correlation relax to their equilibrium $\rho_{\bf q}^{(0)}$ due to the magnon-phonon interaction within the time $\tau_{m}$ [16, 60] via $\partial_{t}{\rho}_{\bf q}|_{\rm rex}=-(\rho_{\bf q}-\rho_{\bf q}^{(0)})/\tau_{m}$. The steady-state solution is obtained by setting $\partial_{t}{\rho}_{\bf q}=0$.
Figure 3(a) plots the real part of magnon correlation density $C_{12}=(1/S_{0}){\rm Re}\sum_{\bf q}\rho_{12}({\bf q})$ injected in an area $S_{0}$ by the spin accumulation $\mu_{s}=\mu_{\uparrow}-\mu_{\downarrow}=0.1$ meV [61, 37, 63, 62] in heavy metals with a lattice constant $a=0.5$ nm and thickness $d=100$ nm [64] at room temperature $T=300$ K. According to Refs. [16, 60], we take $\tau_{m}\approx 1$ ns. The interfacial exchange interactions ${\cal J}_{2}={\cal J}$ and ${\cal J}_{1}={\cal J}\Delta$ can be asymmetric with $\Delta\in[0,1]$, for which we take ${\cal J}=-15$ meV [57, 58, 66, 67, 65]. As resolved in the Brillouin zone in Fig. 3(b), the injected magnon correlation is much larger than the population. Therefore, the magnon population that carries no spin along the Néel vector is only weakly excited owing to spin non-conservation at the interface.
Figure 3: (a) Exchange asymmetry $\Delta$ and magnetic-field dependencies of injected magnon correlation density. (b) Distribution of the injected magnon correlation and population of mode “1” in the Brillouin zone. (c) Injected spin density $S_{z}$ along the Néel vector. (d) Calculated dephasing time $\tau_{\bf q}$ due to the magnon-electron interaction, showing excellent agreement between numerical and analytical calculations.
The spin density $S_{z}$ along the Néel vector is solely carried by magnon correlation. Indeed, as compared by Fig. 3(c) and (a), the injected spin density $S_{z}$ is proportional to $C_{12}$. The net spin injection is suppressed by decreasing the asymmetry $\Delta$ and increasing the magnetic field $H$. This can be attributed to the spin dephasing due to the transfer of magnon spin to the conductor and the intrinsic torque exerted on magnon correlation, which restricts spin injection, as analyzed below.
Focusing on spin dephasing due to magnon-electron coupling, we set $\mu_{\uparrow}=\mu_{\downarrow}$ in Eq. (7). Non-equilibrium magnon correlations now transfer spin angular momentum back to electrons. Disregarding two-magnon excitations $\langle\hat{b}^{\dagger}_{\bf q}\hat{b}^{\dagger}_{-\bf q}\rangle$, the magnon correlation $\rho_{12}({\bf q})$ obeys
${\partial\rho_{12}({\bf q})}/{\partial t}=i\delta\omega_{\bf q}\rho_{12}({\bf q})-\left({1}/{\tau_{\bf q}}+{1}/{\tau_{m}}\right)\rho_{12}({\bf q}).$ (8)
The frequency difference $|\delta\omega_{\bf q}|=|\omega_{1,{\bf q}}-\omega_{2,{\bf q}}|$ that is increased by applied magnetic fields [inset in Fig. 2(d)] causes the oscillation of the correlation. It causes free-induction decay of $S_{z}$ due to the interference of different oscillation frequencies, which is enhanced by the magnetic field. On the other hand,
$\frac{1}{\tau_{\bf q}}=\frac{1}{\hbar}\frac{Sa^{2}}{2d^{2}}C_{q}\sum_{\alpha\beta}\sum_{\zeta=1,2}\hbar\omega_{\zeta,{\bf q}}|\boldsymbol{\sigma}_{\alpha\beta}\cdot\boldsymbol{\chi}^{(\zeta)}_{\bf q}|^{2}$ (9)
is responsible for the pure decay of correlations. Here, $C_{q}=a^{2}dm^{2}/(4\pi\hbar^{4}q)$; the decay rate $1/\tau_{\bf q}$ does not depend on the magnetic field but is suppressed by increasing the exchange asymmetry $\Delta$ as in Fig. 3(d). This is because according to Eq. (6), the coupling constant $\sum_{\alpha\beta}\hbar\omega_{\zeta}({\bf q})|\boldsymbol{\sigma}_{\alpha\beta}\cdot\boldsymbol{\chi}^{(\zeta)}_{\bf q}|^{2}\sim 8JS{\cal J}^{2}(1-\Delta)^{2}$ is suppressed by increasing $\Delta\in[0,1]$. The net dephasing of $S_{z}$ is governed by the joint effect of free-induction decay $\delta\omega_{\bf q}$ and the decay rates $1/\tau_{\bf q}$ and $1/\tau_{m}$.
Gated nonlocal spin transport—The injected magnon correlations carry net spins along the Néel vector and diffuse away from the injector located around $z=0$, resulting in a spin current as depicted in Fig. 4(a). In the steady state, the diffusion equation [50]
$D_{\bf q}\partial^{2}_{z}\rho_{12}(z,{\bf q})-{\rho_{12}(z,{\bf q})}/{\tau_{\rm eff}({\bf q})}=0$ (10)
is governed by the diffusion coefficient $D_{\bf q}=v_{z}^{2}({\bf q})\tau_{\rm eff}({\bf q})$ and the effective dephasing rate $1/\tau_{\rm eff}({\bf q})=1/\tau_{m}-i\delta\omega_{\bf q}$, with $v_{z}({\bf q})=(v_{1,z}({\bf q})+v_{2,z}({\bf q}))/2$ being the group velocity of magnon correlation. With the correlation accumulation $\rho^{(0)}_{12}({\bf q})$ beneath the injector and according to Eq. (4a), Eq. (10) leads to the spin-current density in the region $z>0$ [50]
$J_{z}(z)$ $=\frac{2}{S_{0}}\sum_{\bf q}|v_{z,{\bf q}}|{\cal A}_{\bf q}{\rm Re}\left[\rho^{(0)}_{12}({\bf q})\exp(-\frac{z}{|v_{z,{\bf q}}|\tau_{\rm eff}({\bf q})})\right].$
Figure 4: Nonlocal spin transport tunable by the magnetic fields and metallic gate. (a) Schematic of a nonlocal spin transport device. (b) Diffusive spin current density $J_{z}(z)$ along the $\hat{\bf z}$-direction. Inset: Hanle oscillations at high magnetic field $\mu_{0}H=5$ T. (c) Gated spin transport by normal metals. (d) Spin diffusion length with and without metallic gates, also showing magnetic-field tunability.
Figure 4(b) plots the exponential decay of spin-current densities $J_{z}(z)$ along the transport $\hat{\bf z}$-direction at a small magnetic field $\mu_{0}H=0.5$ T; when the magnetic field is sufficiently strong (e.g., $\mu_{0}H=5$ T), the spin current becomes oscillating as shown in the inset of Fig. 4(b) since in this case $|\delta\omega_{\bf q}|\gg 1/\tau_{m}$, such that $J_{z}(z)$ shows a pronounced Hanle oscillation. This is consistent with experimental observations [31]. As addressed by Eq. (9), the passive metallic gate is a source of spin dephasing. To detect such spin dephasing, we propose to exploit the normal metals to gate the nonlocal spin transport [19] in a device illustrated in Fig. 4(c). The random transfer of spin between normal metals and AFM brakes the spin flow by decreasing the spin diffusion length $\lambda_{s}$ at low magnetic fields through the modified effective dephasing rate $1/\tau_{\rm eff}({\bf q})\rightarrow 1/\tau_{m}+1/\tau_{\bf q}-i\delta\omega_{\bf q}$, as shown in Fig. 4(d). Nevertheless, at higher magnetic fields, the free-induction decay $\delta\omega_{\bf q}$ dominates the spin transport, so the spin flow becomes immune to the metallic gate.
Discussion and conclusion.—The build-up of observable quantum coherence between magnon modes and their transport in thermodynamics is surprising in that no definite phase difference between two modes is promised. This is rooted in a mixing excitation of both modes in a quantum spin-flip process, which creates a magnon correction. The injected magnon correlation oscillates with the frequency difference of two modes, which interfere with each other for different wave vectors. The adjacent metal plays a twofold role: it injects magnon correlation but also causes dephasing of magnon correlation due to their backaction on electron spin dynamics. Our quantum theory establishes a comprehensive understanding of phase-coherent magnon-pair formation in canted AFMs, predicts a practical electrical scheme for manipulating coherent spin signals in thermal magnon transport, and applies to $\alpha$-Fe2O3 [30, 31, 32], FePS3 [68], Cr2O3 [6, 69], and many other noncollinear AFMs.
Acknowledgements.
This work is financially supported by the National Key Research and Development Program of China under Grant No. 2023YFA1406600 and the National Natural Science Foundation of China under Grant No. 12374109.